OpenStudy (anonymous):

What is the equation of the line with an x-intercept of -2 and a y-intercept of 1?

OpenStudy (igreen):

You have two points already. (-2, 0) and (0, 1).

OpenStudy (igreen):

Find the slope, then put it in the form of y = mx + b.

OpenStudy (igreen):

Slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(y_2 = 1\) \(y_1 = 0\) \(x_2 = 0\) \(x_1 = -2\) \(m = \dfrac{1-0}{0--2}\) \(m = \dfrac{1-0}{0+2}\) \(m = \dfrac{1}{2}\) So your slope is 1/2.

OpenStudy (igreen):

y = mx + b Where m = slope and b = y-intercept. You already have the y-intercept and slope so: y = 1/2x + 1

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @__briannnnnaa